# Transition to Petschek Reconnection in Subrelativistic Pair Plasmas: Implications for Particle Acceleration

### v2

@@ 4.2 X-point Injected Particles
Recent work on particle acceleration in relativistic reconnection has demonstrated that the most energetic particles undergo a two-step acceleration process: injection to relativistic energies via direct acceleration at X-points and then strong energization by Fermi/betatron acceleration powered by ideal electric fields in plasmoids (L. Sironi, 2022; S. R. Totorica et al., 2023; S. Gupta et al., 2025). Some particles which fall into the x-point undergo an analogous process in $\sigma=0.3$ reconnection, accelerated by the non-ideal electric field as they undergo Speiser motion (T. W. Speiser, 1965). The result is the establishment of a power-law distribution in the vicinity of the X-point. These particles then are quickly trapped by plasmoids where they ultimately gain most of their energy by Fermi/betatron acceleration as the plasmoid compresses and its magnetic fields intensify.
Let us first consider the process by which the initial power law is formed at the X-point for $\sigma=0.3$ and, specifically, how it differs from that of the $\sigma=10$ case. We follow the procedure of N. Bessho & A. Bhattacharjee (2012) (hereafter BB12) who derived the exit energy spectrum of particles leaving the X-point in the fully-relativistic case. Particles which enter at the center of the diffusion region take the longest to be advected out by the $B_{x}$ field and thus gain the most energy, while those entering near the edges gain the least. To ensure a fair comparison, we compare the primary X-point in the $\sigma=0.3$ case (here the only X-point) with the primary X-point in the $\sigma=10$ case. In BB12 the authors define the edge of the diffusion region as the $E=B$ contour, such that the diffusion region is where the electric field dominates over the magnetic field. This is unsuitable in the $\sigma=0.3$ case since there the $E>B$ zone only covers a small sliver of the diffusion region. We use the modified boundary $E=(v_{A}/c)B$ since this will define the region in which the electric field term in the Lorentz force law will be greater than that of the magnetic field for a particle leaving the diffusion region at the Alfvén speed (Figure 6). As in BB12, we consider a particle as having exited the diffusion region if it passes the maximum or minimum $y$-value of the $E=(v_{A}/c)B$ contour at that point in time. The exit spectrum is thus the time-integrated distribution of exit energies.
@@
Figure 7: Normalized kinetic energy values for tracked simulation electrons upon their departure from the primary diffusion region, summed over time. The dotted line shows the analytical expression for the nonrelativistic exit spectrum (Equation (15)). The humps seen in the curve for $\sigma=0.3$ are caused by a Hall-like $E_{x}$ component noted by M. Swisdak et al. (2008) which results in step-wise acceleration and bunches up the electron distribution; see Appendix C for details. From left to right, the peaks correspond to particles which complete one, two, and three Speiser oscillations before leaving the diffusion region.
If we, like BB12, neglect the Speiser motion in the $x$-direction and repeat their procedure to compute an analytical form for the exit spectrum in the nonrelativistic limit, we arrive at the following result which resembles that of S. V. Bulanov & P. V. Sasorov (1976):
@@
Here $\varepsilon\equiv\sqrt{qE_{z}^{2}/bmc^{2}}$ is a dimensionless parameter quantifying the relative influence of $E_{z}$ on the particle trajectory compared to $B_{y}$, $W^{*}\equiv(qE_{z}^{2}/b)^{-1}W$ is the normalized kinetic energy, and $b\equiv\partial B_{x}/\partial y$, assumed to be constant. Details of the derivation can be found in Appendix B. This fits the data neatly (Figure 7) and may further explain experimental observations of exponential cutoffs in electron energy spectra (W. Fox et al., 2010; B. Na et al., 2023).
@@
Once this power law is established, it is enhanced by acceleration within a plasmoid. Consider the particle shown in Figure 8, a typical case. This electron gains its first substantial amount of energy upon interaction with the X-point. However, most of its ultimate energy gain comes from being confined within a single plasmoid as it compresses and its magnetic fields intensify (F. Guo et al., 2019; H. Hakobyan et al., 2021). This energization can persist so long as the plasmoid continues to accrete more flux and compress adiabatically, heating the particles within. Once the plasmoid is completely immersed within the exhaust, it is deprived of fresh magnetic flux, starving this acceleration mechanism of additional free energy. As a result, the vast majority of plasmoid heating occurs while the plasmoid is still close to the X-point: past $\gtrsim 1000\;c/\omega_{p}$ skin depths, the plasmoids no longer grow and the particles within the plasmoid subsequently no longer gain energy. Thus, unlike in relativistic reconnection, this acceleration pathway has a fixed window of opportunity and cannot generate arbitrarily high-energy particles provided a large enough domain.
@@ 4.3 Exhaust Spectrum
In the large-domain limit only a vanishingly small fraction of particles interact with the X-point at all. The vast majority are simply heated upon interaction with the exhaust. The basic mechanism is pickup acceleration as described in J. F. Drake et al. (2009): when a particle with negligible energy encounters a magnetohydrodynamic flow, the $E$ and $B$ fields accelerate the particle up to speed with the flow. The resultant $E\times B$ cycloid motion imparts a total mean kinetic energy of $m(cE/B)^{2}=mv_{A}^{2}$, equally partitioned between the gyromotion and the guiding-center motion. This acceleration process is responsible for the spectral shoulder in Figure 5, with the mean pickup energy indicated by the dotted line. Once within the flow, these accelerated particles are re-magnetized and undergo mirror motion along the curved field lines, bouncing between opposite boundaries of the exhaust.
The net result of this particle energization is a thermal bulk flow proceeding at the Alfvén speed, with the energy distribution taking Maxwellian form in the comoving frame with $\theta\equiv T/mc^{2}=0.06$. Since the vast majority of the total dissipation occurs in these exhausts and not in X-points/plasmoids, we thus see that while high-$\sigma$ reconnection routinely generates strong power laws in energy, low-$\sigma$ reconnection merely heats the inflows (though a small, energetic contribution within the plasmoids is present). This is consistent with recent work emphasizing the importance of injection at X-points (L. Sironi, 2022; S. R. Totorica et al., 2023; S. Gupta et al., 2025): while the particles which happen to be injected there do reach nonthermal energies, the overwhelming majority which do not will simply be part of the thermal bulk.
@@ 5 Discussion and Conclusions
Secondly, the transition between the plasmoid- and exhaust-dominated reconnection is key for predicting particle acceleration. Although debate remains regarding the relative importance of direct $E$-field acceleration as opposed to Fermi mechanisms (e.g., F. Guo et al., 2025; O. French et al., 2025), our results demonstrate the necessity of plasmoid chains for power law generation. By allowing for stochastic acceleration at multiple X-points and within merging and compressing plasmoids, particles in relativistic reconnection can rapidly form a power-law distribution, avenues that are highly suppressed in an exhaust geometry. This has important effect on understanding whether nonrelativistic electron-ion reconnection is capable of generating strong power laws like those seen in simulations of high-$\sigma$ pair plasmas (e.g., X. Li et al., 2021). Since the injection physics is insensitive to magnetization (Section 4.2), we suggest that the answer will depend on whether the secondary tearing rate is high enough to generate frequent plasmoid mergers. Whether this would be observed in simulations depends on whether the box sizes are large enough to capture such dynamics.
@@
Lastly, our findings provide support for Petschek’s reconnection model. While there has been debate over whether Petschek reconnection can be realized from first-principles simulations (i.e., without ad hoc anomalous resistivity), here we demonstrate a simple such case in which clear, stable laminar Petschek outflows form. While previous work suggested similar conclusions (M. Swisdak et al., 2008; Y.-H. Liu et al., 2012; M. E. Innocenti et al., 2015), ours extends these results in two main ways. Because we use outflow boundary conditions, we can wait for our simulations to reach a steady state, neither a transient nor sensitive to the initial conditions. Also, by simulating much larger domains (thousands of $c/\omega_{p}$), we confirm that plasmoid activity becomes less, not more, impactful as the box size is increased, meaning that these results can be extended to macroscopic system sizes. This bolsters the argument that the anomalous resistivity invoked in MHD simulations and theories of reconnection can indeed result from a fully self-consistent kinetic treatment. Ultimately, our results provide a physical framework for explaining the slow-mode shocks observed in the heliosphere as consequences of Petschek reconnection (e.g., N. K. Walia et al., 2024).
@@
Regarding the localization of the diffusion region, our simulations agree with the suggestion of W. Daughton & H. Karimabadi (2007) that it is the tearing instability which prevents the elongation of the diffusion region by fragmenting the thin current sheet at the X-point should it grow too large. This localization then enables fast Petschek-type reconnection. Hence, although plasmoids are too few in number to generate a plasmoid chain in subrelativistic reconnection, they do play an important role in regulating the structure of the diffusion region.
@@
We leave to further research the extension of these results to electron-ion plasmas and the important question of the role played by three-dimensional effects. Ultimately, the question of whether 3D reconnection proceeds as a dynamic multiple X-point phase or as a predominantly laminar single X-point configuration remains, although flux-rope breakdown (e.g., Q. Zhang et al., 2021) and the drift-kink instability (e.g., F. Bacchini et al., 2025) complicate the simple 2D picture.
@@ Appendix B Derivation of Nonrelativistic Exit Spectrum
This is similar to that found by N. Bessho & A. Bhattacharjee (2012) in the relativistic limit, though there the $W$ dependence in the exponential is $W^{1/2}$ instead of $W^{3/4}$.
@@ Appendix C Hall-like Fields in the Diffusion Region
At the center of the $\sigma=0.3$ reconnection domain lies a tearing-unstable current sheet at the X-point. Given the unity mass ratio, charge separation and the establishment of Hall fields would not be expected. However, strong electrostatic fields comparable in strength to the reconnection electric field are a robust feature of $\sigma=0.3$ systems, appearing in all such simulations conducted (Figure 9). This field structure was noted earlier by M. Swisdak et al. (2008) in similar simulations. Like Hall fields, these electric fields are directed along the $x$-direction. Due to the unity mass ratio, either electric field polarity is possible: for definiteness, we take the ion-like species to be electrons and the electron-like species to be positrons, as develops in Run I. This causes electrons to undergo step-like acceleration as they gain and lose energy to this electric potential over the course of their Speiser orbits, resulting in the discrete bunching of particle energies as seen in Figure 7. It also results (in this polarity) in the electrons possessing more energy than the positrons since they are attracted by the potential. The shapes of the resultant electron and positron energy distributions are nearly identical, with the electron energies simply shifted to the right in the final spectrum. This effect is entirely absent in the $\sigma=10$ case.

### v3

@@ 4.2 X-point Injected Particles
Recent work on particle acceleration in relativistic reconnection has demonstrated that the most energetic particles undergo a two-step acceleration process: injection to relativistic energies via direct acceleration at X-points and then strong energization by Fermi/betatron acceleration powered by ideal electric fields in plasmoids (sironi2022; totorica2023; gupta2025). Some particles which fall into the x-point undergo an analogous process in $\sigma=0.3$ reconnection, accelerated by the non-ideal electric field as they undergo Speiser motion (speiser1965). The result is the establishment of a power-law distribution in the vicinity of the X-point. These particles then are quickly trapped by plasmoids where they ultimately gain most of their energy by Fermi/betatron acceleration as the plasmoid compresses and its magnetic fields intensify.
Let us first consider the process by which the initial power law is formed at the X-point for $\sigma=0.3$ and, specifically, how it differs from that of the $\sigma=10$ case. We follow the procedure of bessho2012 (hereafter BB12) who derived the exit energy spectrum of particles leaving the X-point in the fully-relativistic case. Particles which enter at the center of the diffusion region take the longest to be advected out by the $B_{x}$ field and thus gain the most energy, while those entering near the edges gain the least. To ensure a fair comparison, we compare the primary X-point in the $\sigma=0.3$ case (here the only X-point) with the primary X-point in the $\sigma=10$ case. In BB12 the authors define the edge of the diffusion region as the $E=B$ contour, such that the diffusion region is where the electric field dominates over the magnetic field. This is unsuitable in the $\sigma=0.3$ case since there the $E>B$ zone only covers a small sliver of the diffusion region. We use the modified boundary $E=(v_{A}/c)B$ since this will define the region in which the electric field term in the Lorentz force law will be greater than that of the magnetic field for a particle leaving the diffusion region at the Alfvén speed (Figure 6). As in BB12, we consider a particle as having exited the diffusion region if it passes the maximum or minimum $y$-value of the $E=(v_{A}/c)B$ contour at that point in time. The exit spectrum is thus the time-integrated distribution of exit energies.
@@
Figure 7: Normalized kinetic energy values for tracked simulation electrons upon their departure from the primary diffusion region, summed over time. The dotted line shows the analytical expression for the nonrelativistic exit spectrum (Equation (15)). The humps seen in the curve for $\sigma=0.3$ are caused by a Hall-like $E_{x}$ component noted by swisdak2008 which results in step-wise acceleration and bunches up the electron distribution; see Appendix C for details. From left to right, the peaks correspond to particles which complete one, two, and three Speiser oscillations before leaving the diffusion region.
If we, like BB12, neglect the Speiser motion in the $x$-direction and repeat their procedure to compute an analytical form for the exit spectrum in the nonrelativistic limit, we arrive at the following result which resembles that of bulanov76:
@@
Here $\varepsilon\equiv\sqrt{qE_{z}^{2}/bmc^{2}}$ is a dimensionless parameter quantifying the relative influence of $E_{z}$ on the particle trajectory compared to $B_{y}$, $W^{*}\equiv(qE_{z}^{2}/b)^{-1}W$ is the normalized kinetic energy, and $b\equiv\partial B_{x}/\partial y$, assumed to be constant. Details of the derivation can be found in Appendix B. This fits the data neatly (Figure 7) and may further explain experimental observations of exponential cutoffs in electron energy spectra (fox2010; na2023).
@@
Once this power law is established, it is enhanced by acceleration within a plasmoid. Consider the particle shown in Figure 8, a typical case. This electron gains its first substantial amount of energy upon interaction with the X-point. However, most of its ultimate energy gain comes from being confined within a single plasmoid as it compresses and its magnetic fields intensify (guo2019; hakobyan2021). This energization can persist so long as the plasmoid continues to accrete more flux and compress adiabatically, heating the particles within. Once the plasmoid is completely immersed within the exhaust, it is deprived of fresh magnetic flux, starving this acceleration mechanism of additional free energy. As a result, the vast majority of plasmoid heating occurs while the plasmoid is still close to the X-point: past $\gtrsim 1000\;c/\omega_{p}$ skin depths, the plasmoids no longer grow and the particles within the plasmoid subsequently no longer gain energy. Thus, unlike in relativistic reconnection, this acceleration pathway has a fixed window of opportunity and cannot generate arbitrarily high-energy particles provided a large enough domain.
@@ 4.3 Exhaust Spectrum
In the large-domain limit only a vanishingly small fraction of particles interact with the X-point at all. The vast majority are simply heated upon interaction with the exhaust. The basic mechanism is pickup acceleration as described in drake2009: when a particle with negligible energy encounters a magnetohydrodynamic flow, the $E$ and $B$ fields accelerate the particle up to speed with the flow. The resultant $E\times B$ cycloid motion imparts a total mean kinetic energy of $m(cE/B)^{2}=mv_{A}^{2}$, equally partitioned between the gyromotion and the guiding-center motion. This acceleration process is responsible for the spectral shoulder in Figure 5, with the mean pickup energy indicated by the dotted line. Once within the flow, these accelerated particles are re-magnetized and undergo mirror motion along the curved field lines, bouncing between opposite boundaries of the exhaust.
The net result of this particle energization is a thermal bulk flow proceeding at the Alfvén speed, with the energy distribution taking Maxwellian form in the comoving frame with $\theta\equiv T/mc^{2}=0.06$. Since the vast majority of the total dissipation occurs in these exhausts and not in X-points/plasmoids, we thus see that while high-$\sigma$ reconnection routinely generates strong power laws in energy, low-$\sigma$ reconnection merely heats the inflows (though a small, energetic contribution within the plasmoids is present). This is consistent with recent work emphasizing the importance of injection at X-points (sironi2022; totorica2023; gupta2025): while the particles which happen to be injected there do reach nonthermal energies, the overwhelming majority which do not will simply be part of the thermal bulk.
@@ 5 Discussion and Conclusions
Secondly, the transition between the plasmoid- and exhaust-dominated reconnection is key for predicting particle acceleration. Although debate remains regarding the relative importance of direct $E$-field acceleration as opposed to Fermi mechanisms (e.g., guo2025; french2025), our results demonstrate the necessity of plasmoid chains for power law generation. By allowing for stochastic acceleration at multiple X-points and within merging and compressing plasmoids, particles in relativistic reconnection can rapidly form a power-law distribution, avenues that are highly suppressed in an exhaust geometry. This has important effect on understanding whether nonrelativistic electron-ion reconnection is capable of generating strong power laws like those seen in simulations of high-$\sigma$ pair plasmas (e.g., li2021). Since the injection physics is insensitive to magnetization (Section 4.2), we suggest that the answer will depend on whether the secondary tearing rate is high enough to generate frequent plasmoid mergers. Whether this would be observed in simulations depends on whether the box sizes are large enough to capture such dynamics.
@@
Lastly, our findings provide support for Petschek’s reconnection model. While there has been debate over whether Petschek reconnection can be realized from first-principles simulations (i.e., without ad hoc anomalous resistivity), here we demonstrate a simple such case in which clear, stable laminar Petschek outflows form. While previous work suggested similar conclusions (swisdak2008; liu2012; innocenti2015), ours extends these results in two main ways. Because we use outflow boundary conditions, we can wait for our simulations to reach a steady state, neither a transient nor sensitive to the initial conditions. Also, by simulating much larger domains (thousands of $c/\omega_{p}$), we confirm that plasmoid activity becomes less, not more, impactful as the box size is increased, meaning that these results can be extended to macroscopic system sizes. This bolsters the argument that the anomalous resistivity invoked in MHD simulations and theories of reconnection can indeed result from a fully self-consistent kinetic treatment. Ultimately, our results provide a physical framework for explaining the slow-mode shocks observed in the heliosphere as consequences of Petschek reconnection (e.g., walia2024).
@@
Regarding the localization of the diffusion region, our simulations agree with the suggestion of daughton2007 that it is the tearing instability which prevents the elongation of the diffusion region by fragmenting the thin current sheet at the X-point should it grow too large. This localization then enables fast Petschek-type reconnection. Hence, although plasmoids are too few in number to generate a plasmoid chain in subrelativistic reconnection, they do play an important role in regulating the structure of the diffusion region.
@@
We leave to further research the extension of these results to electron-ion plasmas and the important question of the role played by three-dimensional effects. Ultimately, the question of whether 3D reconnection proceeds as a dynamic multiple X-point phase or as a predominantly laminar single X-point configuration remains, although flux-rope breakdown (e.g., zhang2021) and the drift-kink instability (e.g., bacchini2025) complicate the simple 2D picture.
@@ Appendix B Derivation of Nonrelativistic Exit Spectrum
This is similar to that found by bessho2012 in the relativistic limit, though there the $W$ dependence in the exponential is $W^{1/2}$ instead of $W^{3/4}$.
@@ Appendix C Hall-like Fields in the Diffusion Region
At the center of the $\sigma=0.3$ reconnection domain lies a tearing-unstable current sheet at the X-point. Given the unity mass ratio, charge separation and the establishment of Hall fields would not be expected. However, strong electrostatic fields comparable in strength to the reconnection electric field are a robust feature of $\sigma=0.3$ systems, appearing in all such simulations conducted (Figure 9). This field structure was noted earlier by swisdak2008 in similar simulations. Like Hall fields, these electric fields are directed along the $x$-direction. Due to the unity mass ratio, either electric field polarity is possible: for definiteness, we take the ion-like species to be electrons and the electron-like species to be positrons, as develops in Run I. This causes electrons to undergo step-like acceleration as they gain and lose energy to this electric potential over the course of their Speiser orbits, resulting in the discrete bunching of particle energies as seen in Figure 7. It also results (in this polarity) in the electrons possessing more energy than the positrons since they are attracted by the potential. The shapes of the resultant electron and positron energy distributions are nearly identical, with the electron energies simply shifted to the right in the final spectrum. This effect is entirely absent in the $\sigma=10$ case.